# When Personas Override Payoffs: Role Identity Bias in Multi-Agent LLM Decision-Making

### v2

@@ 2. Related Work
2.1. LLMs in Multi-Agent Systems
Large language models have increasingly been studied as agents in multi-agent systems, with applications spanning negotiation, deliberation, debate, and collaborative problem-solving (Abdelnabi et al., 2024; Bianchi et al., 2024b; Hua et al., 2024; Wang et al., 2025; Liang et al., 2024; Tran et al., 2025; Guo et al., 2024; Xu et al., 2024a; Agashe et al., 2025; Bonabeau, 2002; Epstein and Axtell, 1996). Early work demonstrated that LLMs can sustain multi-turn interactions and coordinate actions through natural language communication while exhibiting emergent collective behaviors (Xu et al., 2024a; Bo et al., 2024). Following studies explored deliberative and collaborative agent frameworks, emphasizing consensus formation and group-level reasoning (Abdelnabi et al., 2024; Bianchi et al., 2024b; Agashe et al., 2025; Bo et al., 2024).
Beyond task completion, multi-agent LLM systems are increasingly used as agent-based simulations to study complex social, economic, and institutional dynamics, reflecting a broader resurgence of agent-based modeling for analyzing emergent behavior in policy and social systems (Bonabeau, 2002; Epstein and Axtell, 1996; Park et al., 2022; Aher et al., 2023; Zheng et al., 2020; Gao et al., 2024).
However, most of this literature focuses on task-oriented collaboration or collective sense-making, rather than strategic settings in which agents must reason about incentives, anticipate others’ actions, and converge to equilibria (Akata et al., 2025; Fan et al., 2023; Herr et al., 2024; Huang et al., 2025; Duan et al., 2024b). As a result, core game-theoretic questions such as payoff optimization, equilibrium selection, and incentive compatibility remain underexplored in multi-agent LLM research.
2.2. LLMs and Game-Theoretic Reasoning
Classical game theory models agents as rational payoff maximizers who compute Nash equilibria under complete information (Nash, 1951). These game-theoretic models have been particularly influential in environmental decision making, where multi-actor coordination failures give rise to socially suboptimal outcomes such as over-extraction and pollution, while emphasizing the biospheric and socio-technical limits of purely objective-driven optimization (Hardin, 1968; Ostrom, 1990; Barrett, 2003; Reyes-Cruz et al., 2024; Domínguez Hernández et al., 2024). A growing body of work has examined whether LLMs approximate such behavior in strategic games, often focusing on simple two-player settings such as Prisoner’s Dilemma, coordination games, or public goods games (Akata et al., 2025; Fan et al., 2023; Herr et al., 2024; Fish et al., 2025). These studies consistently show that LLMs exhibit cooperative or norm-aligned behavior even when defection or self-interested actions are payoff-dominant (Akata et al., 2025; Fan et al., 2023; Herr et al., 2024). Several works suggest that apparent equilibrium behavior in LLMs may arise through semantic or narrative heuristics rather than explicit optimization over incentives (Herr et al., 2024; Macmillan-Scott and Musolesi, 2024; Fish et al., 2025; Bender et al., 2021). When payoffs are implicit, models can sometimes reach Nash equilibria “by accident,” but this behavior is unstable and sensitive to framing. Existing studies rarely isolate the effect of payoff visibility or systematically compare implicit and explicit payoff representations, particularly in multi-agent settings beyond two players (Huang et al., 2025; Duan et al., 2024b; Sun et al., 2025; Duan et al., 2024a). As a result, it remains unclear whether explicit payoff matrices enable principled game-theoretic reasoning or whether LLMs continue to rely on semantic shortcuts even when incentives are fully specified.
2.3. Role Personas and Identity Bias
Role-based prompting and persona conditioning are widely used to steer LLM behavior, with prior work demonstrating effects in persuasion, negotiation, and debate (Abdelnabi et al., 2024; Bianchi et al., 2024b; Hua et al., 2024; Wang et al., 2025; Liang et al., 2024; Neumann et al., 2025). Personas are often introduced to improve realism or alignment with human expectations and are commonly treated as stylistic or surface-level interventions (Park et al., 2023, 2022). However, more recent work has identified a stronger phenomenon commonly referred to as role identity bias, which refers to the tendency of LLMs to adhere to persona-consistent actions even when they conflict with explicit goals, instructions, or incentives, reflecting broader fairness and accountability concerns about how institutional roles and power are encoded in algorithmic systems (Neumann et al., 2025; Macmillan-Scott and Musolesi, 2024; Aher et al., 2023; Rakova and Dobbe, 2023). For example, agents assigned roles associated with profit maximization or authority may favor actions consistent with those identities despite countervailing information (Neumann et al., 2025). Prior studies largely focus on single-agent or two-player interactions, leaving open how role identity bias manifests in larger strategic systems and whether it interferes with equilibrium convergence when explicit payoff information is available (Akata et al., 2025; Herr et al., 2024; Huang et al., 2025). These questions are especially consequential in environmental policy settings, where agents are explicitly cast as stakeholders with conflicting priorities, and where equilibrium selection determines whether coordination or collective-action failure emerges (Ostrom, 1990; Barrett, 2003).
2.4. Information Presentation and Reasoning Fidelity
A considerable amount of literature examines how information presentation affects LLM reasoning across domains, including mathematical problem solving, logical inference, and strategic decision-making, emphasizing that linguistic framing and role description can materially alter behavior even when underlying information is unchanged (Wei et al., 2022; Duan et al., 2024a, b; Herr et al., 2024; Kim et al., 2024; Neumann et al., 2025; Bender et al., 2021; Blodgett et al., 2020). Structured representations, explicit constraints, and intermediate reasoning steps often improve performance, but gains are inconsistent and highly task-dependent (Wei et al., 2022; Duan et al., 2024a, b). Chain-of-thought (CoT) prompting has been shown to improve reasoning accuracy by encouraging intermediate decomposition (Wei et al., 2022). Most CoT research focuses on single-agent tasks, with limited investigation into multi-agent strategic interaction (Duan et al., 2024b, a; Huang et al., 2025). Moreover, recent work cautions that improved surface-level reasoning does not necessarily imply faithful optimization or robust decision-making (Casper et al., 2024; Kim et al., 2024; Weidinger et al., 2022a). The relative importance of payoff visibility versus reasoning scaffolds in achieving equilibrium behavior therefore remains insufficiently understood (Herr et al., 2024; Duan et al., 2024b; Neumann et al., 2025).
2.5. Model and Architectural Effects in Strategic Behavior
Differences in model architecture and training methods have been shown to affect LLM behavior in strategic and social reasoning tasks (Akata et al., 2025; Herr et al., 2024; Huang et al., 2025; Duan et al., 2024b). Model scale alone does not reliably predict performance in strategic settings, and different model families exhibit distinct inductive biases and behavioral tendencies (Akata et al., 2025; Herr et al., 2024; Duan et al., 2024b). Comparative studies suggest that some models rely more heavily on semantic cues and role-based associations, while others appear less sensitive to persona framing (Neumann et al., 2025; Macmillan-Scott and Musolesi, 2024). These findings imply that strategic reasoning in LLMs may be implemented through qualitatively different mechanisms across architectures, with important implications for robustness and deployment in high-stakes settings (Weidinger et al., 2022a; Casper et al., 2024).
Taken together, these strands of work point to an incomplete understanding of strategic behavior in multi-agent LLM systems. Prior work has examined collaboration, game-theoretic reasoning, personas, and information presentation largely in isolation (Guo et al., 2024; Tran et al., 2025; Sun et al., 2025; Akata et al., 2025; Neumann et al., 2025; Kim et al., 2024), but prior existing work has not systematically studied how these factors interact in multi-agent settings with explicit payoff optimization (Sun et al., 2025; Huang et al., 2025; Duan et al., 2024b; Fish et al., 2025). Specifically, it remains unclear whether explicit payoff information enables true game-theoretic reasoning, how role identity bias interferes with equilibrium convergence, and whether these effects vary across model architectures (Fan et al., 2023; Herr et al., 2024; Neumann et al., 2025; Akata et al., 2025). Our work addresses this gap by jointly analyzing role identity, payoff visibility, and reasoning scaffolds in four-agent strategic games, revealing systematic failures of payoff-optimal reasoning and model-dependent patterns of equilibrium selection (Sun et al., 2025; Duan et al., 2024b; Casper et al., 2024).
3. Experimental Design
Our experiment design focuses on distinguishing between identity-driven behavior and payoff-optimal reasoning. This required two key decisions: selecting an appropriate domain that would create a testable tension between role identity bias and strategic payoff maximization, and developing an experimental design that could systematically test how these mechanisms interact in LLM decision-making. We first introduce the domain we chose and explain why it is well-suited for this research, then describe the game structure and experimental design.
3.1. Domain Choice
We chose environmental decision-making as our experimental domain because it creates a natural tension between semantic associations and strategic optimization. Environmental policy involves multiple stakeholders (industry, government, activists, and citizens) with clearly defined but often conflicting interests that map directly to game-theoretic player types. This enables a 4-player game structure that captures real-world dynamics.
Most critically, role-based personas create a conflict between identity-driven behavior and payoff-optimal reasoning. Actions aligned with positive semantic associations from training data may not be payoff-optimal, particularly in economic scenarios where Nash equilibrium may conflict with these associations. In economic scenarios, the Nash equilibrium that maximizes payoffs may involve actions that conflict with these positive associations. This creates a natural experimental setup. If role identity bias dominates, models choose persona-aligned actions, and if game-theoretic reasoning dominates, models choose payoff-optimal actions even when they conflict with role identity. By systematically manipulating persona presence and payoff visibility, we determine whether role identity bias prevents payoff-optimal equilibria and under what conditions payoff-optimal reasoning can be recovered. Our experiments use a 4-player environmental decision-making strategy game where LLM agents play different roles, as shown in Fig. 1:
  1. (1)
    Industrialist (I): chooses between POLLUTE, representing profit-maximizing production with minimal environmental safeguards, and CLEAN, representing investment in environmentally friendly production at higher short-term cost.
  2. (2)
    Government (G): chooses between NOREG, maintaining the regulatory status quo, and REGULATE, imposing environmental regulations that incur enforcement and political costs.
  3. (3)
    Environmental Activist (A): chooses between NOCAMPAIGN, conserving resources and remaining inactive, and CAMPAIGN, mobilizing public pressure through advocacy and protest.
  4. (4)
    Citizen/Consumer (C): chooses between BUY_CHEAP, prioritizing lower-cost goods and short-term affordability, and SUPPORT_GREEN, accepting higher costs to support environmentally sustainable options.
These four roles were chosen to capture the essential stakeholders in environmental policy decisions. The Industrialist represents private industry facing the trade-off between profit and environmental responsibility. The Government represents regulatory authority balancing economic growth with environmental protection. The Environmental Activist represents advocacy groups pushing for environmental action regardless of economic costs. The Citizen represents the general public facing the tension between environmental values and economic constraints. This structure allows us to distinguish between economic actors (Industrialist and Government) and social actors (Activist and Citizen).
3.1.1. Strategy Profiles and Nash Equilibrium
A strategy profile is a 4-tuple specifying each player’s action, for example ‘(Clean, Regulate, Campaign, SupportGreen)‘. Two profiles are particularly important:
  • Green Transition: ‘(Clean, Regulate, Campaign, SupportGreen)‘ - All players choose environmentally-friendly actions
  • Tragedy of Commons: ‘(Pollute, NoReg, NoCampaign, BuyCheap)‘ - All players prioritize self-interest over environmental concerns
After all agents make their decisions simultaneously, we compute whether the resulting strategy profile is a Nash equilibrium. A strategy profile is a Nash equilibrium if no player can improve their payoff by unilaterally changing their action, given the payoff structure and other players’ actions. If models are using game-theoretic reasoning, they should converge to Nash equilibria. If role identity bias dominates, they may choose profiles that align with their respective persona’s expectations but are not Nash equilibria.
The payoff structure determines which profiles are Nash equilibria. For example, if payoffs are structured so that Green Transition gives the highest payoffs to all players, then Green Transition is a Nash equilibrium. If payoffs are modified so that polluting is more profitable, then Tragedy of Commons might become the Nash equilibrium instead.
3.1.2. Scenarios
We tested 53 scenarios across three scenario types, all designed with payoff-coupled narratives where the scenario description aligns with the underlying payoff structure. All scenario narratives were generated using GPT-5-mini [^1] Azure OpenAI GPT-5-mini (model version 2025-08-07) and subsequently reviewed and filtered to ensure consistency with the intended payoff structure and equilibrium class [^2] Model-generated scenarios were used to reduce author-specific framing bias and to enable systematic variation across narrative surface forms while holding the underlying game structure fixed. Scenarios were retained only if the narrative description unambiguously aligned with the corresponding payoff incentives, and no model was evaluated on scenarios it had previously generated.. Models are exposed to each scenario in two versions: one where payoffs are implicit and must be inferred from the narrative description, and one where explicit payoff matrices are provided. Scenarios are categorized into three subtypes  [^3] Example from the scenario set and a complete list of scenario subtypes are included in Appendix A.7.:
  • Economic/Tragedy Scenarios: These scenarios describe economic crises where the payoff structure makes Tragedy of Commons the dominant Nash equilibrium.
  • Environmental/Green Scenarios: These scenarios describe environmental emergencies where the payoff structure makes Green Transition the dominant Nash equilibrium.
  • Mixed/Complex Scenarios: These scenarios involve situations where economic and environmental considerations are intertwined, creating complex payoff structures that may not clearly favor either Green Transition or Tragedy of Commons as the dominant equilibrium.
The scenario set is intentionally imbalanced across these two classes. Environmental scenarios are overrepresented because they reflect a primary real-world domain in which role-based multi-agent LLM systems are actively deployed for policy analysis, sustainability planning, and climate governance. Economic Tragedy-of-the-Commons scenarios serve as contrast cases, enabling controlled comparison between settings where payoff-optimal actions conflict with positive semantic associations and settings where payoff-optimal actions align with them. This design allows us to test how models balance identity-driven behavior and payoff-optimal reasoning, whether they can infer payoffs from narrative descriptions, whether explicit payoff information enables payoff-optimal reasoning, and how role identity bias interacts with payoff visibility to affect equilibrium selection.
Table 1. Overall Nash Equilibrium Rates: With Personas vs Without Personas. This table compares overall Nash equilibrium rates across all scenarios (economic + environmental) between conditions with and without personas. The “Difference” column shows the change in percentage points when personas are removed (negative values indicate personas help, positive values indicate personas hurt).
Model Payoffs With Persona Without Persona Difference
Llama-8B Hidden 77.4% (205/265) 75.5% (200/265) -1.9 pp
Llama-8B Visible 77.0% (204/265) 72.1% (191/265) -4.9 pp
Mistral-7B Hidden 77.4% (205/265) 75.8% (201/265) -1.6 pp
Mistral-7B Visible 69.8% (185/265) 38.9% (103/265) -30.9 pp
Qwen-32B Hidden 69.1% (183/265) 74.3% (197/265) +5.2 pp
Qwen-32B Visible 57.7% (153/265) 33.2% (88/265) -24.5 pp
Qwen-7B Hidden 75.5% (200/265) 75.5% (200/265) 0.0 pp
Qwen-7B Visible 75.5% (200/265) 55.5% (147/265) -20.0 pp
3.2. Chain-of-Thought Reasoning Analysis
We analyze the chain-of-thought reasoning generated by models during action selection to understand how role identity bias manifests in decision-making processes. For each agent’s action selection, we extract the rationale field containing the model’s reasoning and analyze keyword patterns that distinguish different reasoning mechanisms (see Appendix C for keyword categories). We compare keyword frequency distributions across experimental conditions to identify how personas and payoff visibility affect reasoning style. From this analysis, we infer whether models are using game-theoretic reasoning (direct payoff optimization) or identity-driven reasoning (role-aligned decision-making). This allows us to understand the mechanisms behind observed Nash equilibrium patterns and how role identity bias shifts reasoning from payoff-optimal to identity-aligned strategies. While we reference key CoT findings in the main text to support our quantitative results, we do not analyze these results in detail here. A comprehensive analysis of keyword frequency patterns, reasoning shifts, and detailed visualizations are presented in the Appendix (Figures 5-8).
3.3. Models and Experimental Conditions
Our experimental design follows a $\times2$ factorial structure, crossing role identity and payoff visibility, and tests 53 environmental decision-making scenarios in a 4-agent simultaneous-move game. We use four instruction-tuned, open-source LLMs: Qwen2.5-7B (Qwen et al., 2025), Qwen2.5-32B (Qwen et al., 2025), Llama-3.1-8B (Grattafiori et al., 2024), and Mistral-7B (Jiang et al., 2023). For each model, condition, and scenario, a single model instance takes on all four agent roles (Industrialist, Government, Environmental Activist, Citizen) and is prompted independently for each role. Agents select actions in a fixed order [^4] This order is fixed (not randomized) to ensure reproducibility across experiments, and each configuration is repeated five times.
We compare two persona conditions: with personas (role-specific descriptions reflecting stakeholder motivations) and without personas (agents instructed only on Nash equilibrium reasoning). We also compare two payoff conditions: hidden (agents infer incentives from scenario descriptions) and visible (agents receive explicit payoff matrices). All experiments use chain-of-thought reasoning instructions. We evaluate outcomes along two dimensions: whether the resulting action profile constitutes a Nash equilibrium and which canonical equilibrium is selected (Green Transition or Tragedy of Commons).

### v3

@@ 2. Related Work
Our work situates at the intersection of several active research areas: the use of LLMs as multi-agent systems in simulations, game theoretic reasoning in LLMs, the behavioral effects of role-based personas, and sensitivity of LLM behavior to prompt design and model choices.
2.1. LLMs as Multi-Agent Systems and Simulation Substrates
Large Language Models (LLMs) are increasingly studied as agents in multi-agent settings. Interactive negotiation games show that LLM agents can sustain multi-turn strategic interactions (Abdelnabi et al., 2024), though controlled benchmarks also reveal systematic irrationalities in their behavior (Bianchi et al., 2024). Structured debate among multiple agents can improve reasoning quality by eliciting diverse intermediate perspectives (Liang et al., 2024). A complementary line of research treats LLM based multi-agent systems as simulation environments for human or institutional behavior. Populations of LLM agents can reproduce patterns observed in human-subject studies (Aher et al., 2023), simulate social environments with persistent memory and behavioral continuity (Park et al., 2022, 2023), generate emergent policy-relevant dynamics at macroeconomic scale (Li et al., 2024), and operate within formalized simulation frameworks that emphasize the role of design choices in shaping emergent behavior (Vezhnevets et al., 2023). These developments connect LLM-based agent systems to the broader tradition of agent-based modeling, where foundational work emphasizes that simulation outputs depend critically on modeling assumptions and representational choices (Bonabeau, 2002; Epstein and Axtell, 1996).
Recent surveys document the rapid growth of LLM-based multi-agent systems, while also noting that much of this literature emphasizes task performance rather than principled modeling of incentives and representations (Guo et al., 2024; Gao et al., 2024; Tran et al., 2025). What remains underexplored is how representational design choices, such as role specification or payoff encoding, shape system-level outcomes in these simulation settings. One natural starting point is to ask whether these agents reason strategically over incentives at all.
2.2. Game-Theoretic Reasoning in LLMs
Early evidence suggests that LLMs can perform nontrivial reasoning about other agents in strategic settings, though such capabilities depend strongly on prompting and problem formulation (Gandhi et al., 2023). Systematic evaluations across canonical games confirm this fragility: models often deviate from rational best-response behavior  (Fan et al., 2023), cooperate more than humans in Prisoner’s Dilemma settings (Fontana et al., 2025), and exhibit unstable strategies in repeated play (Akata et al., 2025). In strategic pricing environments, LLM agents can even exhibit collusive behavior, raising concerns about emergent coordination (Fish et al., 2025).
Benchmark driven work further quantifies these limitations. GTBench reveals substantial weakness in LLM strategic reasoning across structured tasks (Duan et al., 2024b), while recursive look-ahead prompting can improve performance, although it does not eliminate failures (Duan et al., 2024a). More recent work reframes these behaviors through the lens of bounded rationality: both humans and LLMs deviate from Nash equilibrium in ways consistent with heuristic reasoning (Zheng et al., 2025), and LLM decisions often align more closely with behavioral game-theoretic patterns than with strict payoff maximization (Jia et al., 2025).
Across these studies, a consistent picture emerges: LLMs often rely on semantic and heuristic reasoning rather than explicit incentive optimization. Most prior evaluations, however, focus on two-player settings and do not isolate the role of payoff representation. One factor that may contribute, but has received limited attention, is the role-based identity framing used to condition agent behavior.
2.3. Role Personas and Identity Effects
Role prompting is widely used to shape LLM behavior, but a growing body of work suggests that such framing choices can introduce systematic biases. LLMs encode cognitive and social identity biases that affect outputs across a range of contexts  (Macmillan-Scott and Musolesi, 2024; Hu et al., 2025), and system-level prompts can themselves function as structural sources of bias (Neumann et al., 2025b). In simulation settings, richer persona modeling affects behavioral fidelity and perceived realism (Merrill and Srivastava, 2025). This suggests that personas are not merely stylistic devices but active shapers of agent behavior. Yet whether such identity effects interfere with strategic reasoning in multi-agent environments, where equilibrium selection depends on joint behavior, has not been directly examined. A related question is whether these effects depend on how incentives are presented and whether they manifest differently across model architectures.
2.4. Prompt Sensitivity and Model Variation
Prompt design and information presentation can substantially alter LLM behavior. Chain-of-thought prompting can improve reasoning by eliciting intermediate steps (Wei et al., 2022). However, these improvements are often inconsistent and depend on task structure and prompt formulation (Creswell and Shanahan, 2023; Cobbe et al., 2021).
In strategic settings, prompt sensitivity is particularly consequential. LLM decisions can depend more on contextual framing than on the underlying game structure (Lorè and Heydari, 2024). Models exhibit systematic biases in strategic games depending on how information is presented (Herr et al., 2024). Explicit reasoning scaffolds can improve strategic performance (Duan et al., 2024a), but observed model behavior remains highly sensitive to interface and evaluation conditions, complicating interpretation (Casper et al., 2024). More broadly, LLM outputs are strongly shaped by surface form and statistical patterning rather than grounded understanding (Bender et al., 2021), and these sensitivities pose risks when model behavior is misinterpreted or over-trusted (Weidinger et al., 2022).
These effects are not uniform across models. Strategic behavior varies substantially across model families and is not a monotonic function of scale  (Duan et al., 2024b; Huang et al., 2025). Different models adopt different strategies in repeated interactions (Akata et al., 2025) and vary in their susceptibility to framing effects (Herr et al., 2024). This means that the impact of any representational choice may itself be model-dependent, complicating claims about LLM strategic behavior as a general phenomenon.
The preceding sections identify four factors that shape LLM behavior in strategic settings: the multi-agent simulation context, the limits of game-theoretic reasoning, the influence of role-based personas, and the sensitivity to prompt design and model choice. Prior work has studied these factors largely in isolation. What remains missing is a unified analysis of how they interact in multi-agent strategic systems.
In this work, we jointly study role personas, payoff visibility, and model architecture in four-agent strategic games, with supplementary analyses testing robustness to persona prompt formulation. We show that persona conditioning can override explicit incentives, preventing equilibrium attainment even when payoffs are fully specified. Furthermore, we demonstrate that the effect of payoff information depends critically on persona presence and varies across model architectures. These results establish that representational choices are not merely implementation details, but can determine whether a system behaves more like a payoff-driven strategic reasoner or an identity-driven simulator, with direct implications for governance and deployment.
3. Methodology
We study whether LLM agents behave as payoff-sensitive strategic actors or as identity-driven role followers in multi-agent decision settings. Our goal is to understand how representational choices, specifically persona descriptions and payoff presentation, affect strategic behavior.
To study this, we construct a controlled multi-agent decision setting based on an environmental policy game. The game models a simplified governance scenario involving four stakeholders industry, government, activists, and consumers, each making a binary decision that affects both individual outcomes and collective welfare. Depending on how incentives are structured, these interactions can lead either to coordinated pro-environmental outcomes (Green Transition) or to individually beneficial but collectively harmful outcomes (Tragedy of the Commons).
Within this setting, we systematically vary how the decision problem is presented to LLM agents. Specifically, we test whether agents are given role-based personas that describe stakeholder motivations, and whether payoff information is explicitly shown or must be inferred from the narrative. For each configuration, we simulate decisions for all four roles using the same underlying model and analyze the resulting joint outcome.
Fig. 1 shows our overall setup. This allows us to examine whether agents follow explicit incentives or instead default to role-consistent behavior, and how this depends on model family and scenario structure. We now formalize the game and experimental design.
3.1. Experimental Design
We use a $2\times 2$ factorial design to vary how agents are framed and how incentives are presented.
Persona condition In the persona setting, each role is given a short description of its stakeholder identity and motivations. In the no-persona condition, this description is removed and agents are instructed only to act strategically.
Structural role labels (Industrialist, Government, Activist, and Citizen) are retained in both conditions in order to preserve the game definition while isolating the effect of explicit identity framing.
Payoff visibility In hidden-payoff scenarios, agents see only the narrative and must infer incentives. In visible-payoff scenarios, the full payoff table is provided explicitly.
Combining these two factors yields four experimental conditions:
$\{\texttt{persona},\texttt{no-persona}\}\times\{\texttt{hidden},\texttt{visible}\}.$
To evaluate the robustness of persona effects to prompt formulation, we additionally introduce controlled variations within the persona condition. These variants preserve the same game structure, payoff specification, and action space, but systematically alter how agent identities are expressed, including differences in wording, strength of role framing, explicit prioritization of payoffs, and surface-level role labels. This allows us to isolate whether observed behaviors depend on a specific prompt formulation or reflect a more general effect of persona-based representation. Details of the specific prompt variants are provided in Appendix F.
This design enables us to distinguish between identity-driven and payoff-driven behavior. If agents act as payoff maximizers, they should converge to Nash equilibria when incentives are explicit, regardless of persona framing. Conversely, if role identity dominates, behavior should remain aligned with persona-consistent actions even when this conflicts with payoff-optimal strategies. The prompt variations further allow us to test whether these effects are stable under changes in representation, or sensitive to specific formulations of role descriptions.
3.2. Interaction Protocol
For each model, scenario, and condition, we simulate a single round of simulation play. Each of the four roles is prompted independently using the same underlying model instance. The resulting actions are combine into joint strategy profile. Agents select actions in a fixed order across all experiments to ensure reproducibility as no agent observe the choices of others. This ordering does not affect the strategic structure of the game. Each model–scenario–condition combination is repeated five times to account for stochasticity in model outputs.
3.3. Domain and Game Design
We use an environmental policy setting because it captures real-world governance scenarios where multiple stakeholders with conflicting objectives must coordinate under shared constraints. Examples include climate policy, regulation, and resource management, where decisions depend on both incentives and stakeholder roles. This setting creates a natural tension between strategic incentives (captured by payoffs) and role semantics (captured by personas). For instance, an industrial actor may be incentivized to pollute under certain payoffs, while a regulator or activist may be expected to favor environmentally beneficial outcomes. This allows us to test whether LLM agents follow explicit incentives or default to role-consistent behavior.
We define the set of players in the game as:
$N=\{I,G,A,C\}$
corresponding to an Industrialist, Government, Environmental Activist, and Citizen.
Each player $i\in N$ has a binary action space corresponding to a policy-relevant choice:
$A_{I}=\{\texttt{POLLUTE},\texttt{CLEAN}\},A_{G}=\{\texttt{NOREG},\texttt{REGULATE}\},$
$A_{A}=\{\texttt{NOCAMPAIGN},\texttt{CAMPAIGN}\},A_{C}=\{\texttt{BUY\_CHEAP},\texttt{SUPPORT\_GREEN}\}$
A joint strategy profile represents the actions chosen by all four players in a single round of the game:
$a=(a_{I},a_{G},a_{A},a_{C})\in A_{I}\times A_{G}\times A_{A}\times A_{C}$
where each component $a_{i}\in A_{i}$ is the action chosen by player $i$. Since each of the four players has two possible actions, the game contains $2^{4}=16$ possible pure action profiles.
Each scenario defines a payoff function for each player, assigning a real-valued utility to every joint action profile:
$u_{i}(a):A_{I}\times A_{G}\times A_{A}\times A_{C}\rightarrow\mathbb{R},\quad i\in N$
We focus on two canonical profiles that capture the main coordination structures of interest:
$a_{green}=(\texttt{CLEAN},\texttt{REGULATE},\texttt{CAMPAIGN},\texttt{SUPPORT\_GREEN})$
$a_{tragedy}=(\texttt{POLLUTE},\texttt{NOREG},\texttt{NOCAMPAIGN},\texttt{BUY\_CHEAP})$
We refer to these as the Green Transition and the Tragedy of the Commons, representing coordinated pro-environmental action and individually beneficial but collectively harmful behavior, respectively.
In the visible-payoff condition, agents receive a complete payoff table specifying outcomes for all 16 joint action profiles. Each entry provides a 4-tuple of payoffs corresponding to the Industrialist, Government, Activist, and Citizen, allowing agents to directly evaluate the consequences of different strategies. In the hidden-payoff condition, agents must instead infer incentives from the narrative description.
3.4. Equilibrium Definition and Outcome Measures
We define equilibrium using the standard notion of Nash equilibrium, which captures stability under unilateral deviations.
Intuitively, this means that no agent can improve their payoff by unilaterally changing their own action while all other players keep their actions fixed. Equivalently, each player’s action is a best response to the actions of the others, and the resulting joint outcome is stable under individual deviations.
A profile $a^{*}$ is a Nash equilibrium if
$u_{i}(a_{i}^{*},a_{-i}^{*})\geq u_{i}(a_{i},a_{-i}^{*})\quad\forall i\in N,\forall a_{i}\in A_{i}.$
We use Nash equilibrium as an operational diagnostic of incentive-consistent strategic behavior: if agents reason over payoffs, their joint actions should satisfy equilibrium conditions. Importantly, we do not treat Nash equilibrium as a normative objective, but as a tool to measure whether agent behavior aligns with the incentive structure of the game.
For each joint strategy profile $a$, we compute two outcome measures.
Table 1. Representative examples of scenario types used in our experiments, illustrating how narrative framing aligns with underlying payoff structures. Economic scenarios encode payoff–semantic conflict where individually optimal actions lead to socially undesirable outcomes (Tragedy of the Commons). Environmental scenarios encode payoff–semantic alignment where coordinated pro-environmental actions are both socially and individually optimal (Green Transition).
Type Example Scenario
A table presenting three types of experimental scenarios used in the study: Economic and Environmental. Each row contains a detailed narrative description illustrating how incentives and constraints are structured. Economic scenarios depict situations where polluting production is the only viable or most profitable option, creating a conflict between individual incentives and socially desirable outcomes. Environmental scenarios describe conditions where policy mechanisms such as carbon taxes and subsidies make clean production both profitable and aligned with social goals. The table highlights how narrative framing encodes payoff structures that agents must reason about in the experiments.
Nash attainment.
This is a binary indicator of whether the realized outcome is an equilibrium:
$Nash(a)=\begin{cases}1&a\in NE\\ 0&\text{otherwise}\end{cases}$
Equilibrium type.
For outcomes that are Nash equilibria, we additionally classify which kind of equilibrium the model reached.
$\text{Type}(a)=\begin{cases}\textsc{Green Transition}&\text{if }a=a_{\text{green}}\\ \textsc{Tragedy of Commons}&\text{if }a=a_{\text{tragedy}}\\ \textsc{Other}&\text{otherwise}\end{cases}$
If $a\in NE$, we classify whether it corresponds to Green Transition, Tragedy of the Commons, or another equilibrium. Formally, Green-dominant scenarios correspond to $a_{green}\in NE$, and Tragedy-dominant scenarios correspond to $a_{tragedy}\in NE$.
Together, these measures allow us to distinguish between two different phenomena: (i) failure to reach equilibrium at all, and (ii) systematic biases in equilibrium selection when equilibrium is achieved.
3.5. Scenario Construction
We construct 53 scenarios by specifying payoff tables over the 16 joint action profiles and pairing them with natural-language narratives describing the policy context. Scenarios are generated by first choosing a target equilibrium structure, constructing payoffs consistent with that structure, and then generating narratives aligned with the incentives. We manually filter scenarios to ensure that the narrative meaningfully reflects the underlying payoff structure.
To organize the evaluation, we group scenarios according to which equilibrium structure they instantiate:
  • Green-dominant: $a_{\text{green}}\in NE$ and $a_{\text{tragedy}}\notin NE$. Environmentally friendly Green Transition is the stable outcome among two canonical profiles, while Tragedy is not.
  • Tragedy-dominant: $a_{\text{tragedy}}\in NE$ and $a_{\text{green}}\notin NE$. Individually beneficial but collective harmful behavior, Tragedy of the Commons is the stable outcome, while Green Transition is not.
Table 1 shows an example of each type of scenario, and Table 2 in the appendix shows the scenario subtypes. Green-dominant scenarios are those in which coordinated environmentally favorable action forms a stable equilibrium. Tragedy-dominant scenarios are those in which individually beneficial but collectively harmful behavior is strategically stable.
3.6. Models
We evaluate four instruction-tuned open-source models: Qwen2.5-7B and Qwen2.5-32B (Qwen et al., 2025), Llama-3.1-8B (Grattafiori et al., 2024), and Mistral-7B (Jiang et al., 2023).
These models are selected to capture both within-family and across-family variation. The Qwen models provide a limited scaling comparison within the same family, while Llama and Mistral introduce variation in model family and training recipe. All four are instruction-tuned, open-source models that are representative of current LLMs used in reasoning and multi-agent settings. This setup allows us to test whether persona and payoff effects are consistent across models or depend on model-specific inductive biases.
Overall, the experimental pipeline proceeds as follows: we fix a game and scenario, vary persona and payoff conditions, generate independent decisions for each role using a given model, and evaluate the resulting joint outcome in terms of equilibrium attainment and type. By holding the underlying game constant while varying representation, this design isolates whether LLM agents behave as payoff-driven strategic reasoners or as identity-conditioned actors.